OpenStudy (biue):

Will fan and medal! 5 ( n - 1/10 ) = 1/2 Solve

OpenStudy (texaschic101):

5(n - `1/10) = 1/2 -- distribute through the parenthesis 5n - 1/2 = 1/2 -- add 1/2 to both sides 5n = 1/2 + 1/2 5n = 1 n = 1/5 check.. 5(n - 1/10) = 1/2 5(1/5 - 1/10) = 1/2 5(2/10 - 1/10) = 1/2 5(1/10) = 1/2 5/10 = 1/2 1/2 = 1/2 (correct) so.... n = 1/5 any questions ?